OpenStudy (anonymous):

The sum of two numbers is 15, and their difference is 3, what are the two numbers?

OpenStudy (anonymous):

14 and 5

OpenStudy (anonymous):

I edited the question. Please help!

OpenStudy (anonymous):

9 & 6

OpenStudy (anonymous):

Thanks. How did you get that please?

OpenStudy (anonymous):

Write it out as two equations: a + b = 15 a - b = 3 Solve one of them for either a or b (it doesn't matter which equation or which variable): a = 3 + b Plug that into the other equation and solve: (3 + b) + b = 15 3 + 2b = 15 2b = 12 b = 6 Now you can use either equation to solve for a: a - 6 = 3 a = 9 So your answer is 6 and 9 (plug them into both equations to double check).

OpenStudy (anonymous):

oh ok cool! Thanks